The density of a gas at SATP is 2.50g/L. Determine the molecular mass of the gas
1
Expert's answer
2020-06-05T05:53:44-0400

According to the ideal gas law, the equation of state is:

"pV=nRT" ,

where "p" is the pressure, "V" is the volume, "n" is the number of the moles, "R" is the ideal gas constant (8.314 J mol-1 K-1) and "T" is the temperature.

The mass of the gas "m" is the product of the number of the moles and the molar mass "M":

"m = n\u00b7M" .

Joining the two equations above:

"pV = \frac{m}{M}RT"

The density "d" of a gas is the ratio of the mass to the volume:

"d =\frac{m}{V} = \frac{pM}{RT}" .

Therefore, the molar mass of the gas is:

"M = \frac{d}{p}RT" .

Standard conditions for temperature and pressure (SATP) are 105 Pa and 273.15 K. Finally:

"M= \frac{2.50·10^3(\text{g/m}^3)}{10^5(\text{Pa})}·8.314(\text{J/(mol K)})·273.15(\text{K})"

"M = 56.8 \text{ g/mol}" .

Answer: the molar mass of the gas is 56.8 g/mol